# Towards Safe and Robust Autonomous Vehicle Platooning: A Self-Organizing Cooperative Control Framework

### v1

@@ IV-B Hardware-in-loop experiment
Figure 10: Overall Architecture of the HIL Experimental Platform
Figure 11: Schematic representation of case 1 results in HIL experiments
Figure 12: Schematic representation of case 2 results in HIL experiments
@@ IV-B1 HIL experimental platform construction
In the Hardware-in-the-Loop (HIL) experiments, we constructe an advanced simulation platform integrating a domain controller with Carla[53], a high-fidelity driving simulator, to validate our proposed model in a controlled yet realistic environment. The experimental highway scenario is created using Roadrunner, with varied background traffic flows generated using Highway-env.
TABLE III: Specifications of the domain controller
@@
The platform consists of the domain controller, Carla simulation environment, Roadrunner, Highway-env, and PID controllers. The domain controller, whose parameters are shown in Table III, processes real-time traffic and road information from Carla. Carla provides detailed road conditions and vehicle dynamics, which are translated into vehicle movements via PID controllers based on high-level action commands. Roadrunner generates complex highway scenarios, while Highway-env produces traffic with varied driving styles to test the algorithm’s robustness.
@@ IV-B2 Case analysis
This section examines the performance of our model on the HIL simulation platform in real-time autonomous platooning scenarios. The results indicate that our model can effectively perform perception, decision-making, and control, ensuring safe and robust autonomous platooning. We analyze two typical scenarios involving significant static and dynamic interferences to illustrate the model’s strategies under these conditions. Demonstration videos are available on our website  [^1] Hardware-in-the-Loop Experimental Validation Video Weblink.
In the first scenario, the platoon encounters a severe static obstruction: a multi-vehicle collision blocking multiple lanes. Traditional CACC strategies cannot manage coordinated deceleration due to inadequate braking distances, and CLCC strategies are hindered by vehicles behind the accident on the right. Our model resolves this complex situation through adaptive decision-making, avoiding secondary collisions by allowing the first vehicle to initiate a lane change and the other vehicles to follow once adequate space is available.
As shown in Figure 11, the accident is detected at step 30. The first vehicle moves to the rightmost lane, completing the maneuver by step 105. By step 160, the remaining vehicles have sufficient space to change lanes, avoiding the collision zone, and by step 220, all vehicles have successfully moved. The LQR controller then stabilizes the platoon with reduced headway distances, demonstrating the model’s flexibility and efficiency in navigating static obstructions.
In the second scenario, the platoon faces a dynamic obstruction: the leading vehicle decelerates suddenly at 5 m/s² due to a breakdown. Traditional strategies would struggle to avoid collisions due to short following distances and the presence of background traffic. Our model, however, initiates a staggered lane change, as illustrated in Figure 12. The first vehicle moves left at step 20, allowing the second to move right. The third vehicle, facing traffic in the right lane, moves left. By step 70, all vehicles have repositioned and resumed normal speeds. The rightmost lane, offering higher efficiency, sees gradual reformation of the platoon, with all vehicles settled by step 455.
These scenarios demonstrate the model’s capability to handle sudden obstacles by dynamically adjusting the platoon’s formation and leveraging available lane space to maintain safety and efficiency. The adaptive lane-changing strategy ensures minimal disruption and optimal traffic flow, showcasing the model’s robustness and flexibility in complex driving conditions.
@@ V Conclusion
In the complex environment of mixed human and autonomous driving, achieving robust and safe self-organizing decision-making and control for autonomous vehicle platooning remains a significant challenge for commercial deployment. This paper presents a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework for autonomous vehicle platoons. This framework enables adaptive decision-making in various complex scenarios by flexibly switching between reinforcement learning and model control methods, integrating traffic priors and optimal control techniques. Our model ensures safe and stable operation even under severe dynamic and static disturbances. Extensive simulations and HIL tests have demonstrated the effectiveness of the proposed model. The results show that our model significantly outperforms traditional platooning strategies such as CACC and CLCC in terms of safety, efficiency and robustness, demonstrating its potential for real-world deployment.
Overall, this research contributes to the advancement of autonomous vehicle platooning by providing a model that balances safety, efficiency, and robustness. Future work will focus on optimizing the model’s performance, exploring its scalability to larger platoons, and conducting real-world testing to confirm its practical benefits and limitations.

### v2

@@ IV-B Hardware-in-loop experiment
Figure 8: Illustration of the HIL platform architecture for autonomous vehicle platooning, which integrates a CARLA simulation environment, RoadRunner for road scenes and traffic flow, and domin controllers for decision-making and control.
Figure 9: Illustration of the adaptive response to static obstruction in autonomous vehicle platooning, which decipts a scenario where the platoon encounters a multi-vehicle collision, necessitating coordinated lane changes to navigate the obstruction safely.
Figure 10: Illustration of the adaptive response to dynamic obstruction in autonomous vehicle platooning, which depicts a scenario where the platoon leader’s sudden deceleration triggers a staggered lane-change maneuver among following vehicles to prevent collisions and maintain traffic flow.
@@ IV-B1 HIL experimental platform construction
In the HIL experiments, an advanced simulation platform is constructed by integrating a domain controller with Carla [37], a high-fidelity driving simulator, to validate the proposed model in a controlled yet realistic environment. The experimental highway scenario is created using Roadrunner, with varied background traffic flows generated by Highway-env.
TABLE IV: Specifications of the domain controller
@@
The platform consists of the domain controller, Carla simulation environment, Roadrunner, Highway-env, and PID controllers. The domain controller, whose parameters are shown in Table IV, processes real-time traffic and road information from Carla. Carla provides detailed road conditions and vehicle dynamics, which are translated into vehicle movements via PID controllers based on high-level action commands. Roadrunner generates complex highway scenarios, while Highway-env produces traffic with varied driving styles to test the algorithm’s robustness.
@@ IV-B2 Case analysis
This section examines the performance of our model on the HIL simulation platform in real-time autonomous platooning scenarios. The results indicate that our model can effectively perform perception and decision-making, ensuring safe and robust platooning. Two typical scenarios with significant static and dynamic interferences are analyzed to illustrate the model’s strategies. Demonstration videos are available on our website  [^1] Hardware-in-the-Loop Experimental Validation Video Weblink.
In the first scenario, the platoon encounters a multi-vehicle collision blocking several lanes. Traditional strategies struggle with safe deceleration and coordination, but our model enables the lead vehicle to initiate a lane change, followed by others when safe. As shown in Figure 9, the lead vehicle moves at step 30, with all vehicles clearing the obstruction by step 220. The LQR controller stabilizes the platoon with reduced headway, illustrating flexibility in static obstructions. In the second scenario, the lead vehicle suddenly decelerates at 5 m/s² due to a breakdown. Traditional models risk collision due to short distances, but ours initiates a staggered lane change (see Figure 10). By step 70, all vehicles are safely repositioned, and the platoon reforms in the right lane by step 455, maintaining safety and flow.
These scenarios demonstrate the model’s capability to handle sudden obstacles by dynamically adjusting the platoon’s formation and leveraging available lane space to maintain safety and efficiency.
@@ V Conclusion
In the evolving field of autonomous driving, ensuring safe and robust self-organizing decision-making for vehicle platooning, especially in mixed traffic, presents significant challenges. This paper introduces TriCoD, a Twin-World Safety-Enhanced Data-Model-Knowledge Triple-Driven Adaptive Autonomous Vehicle Platooning Cooperative Decision-Making Framework. By integrating reinforcement learning with model control techniques, TriCoD enables adaptive decision-making across diverse, complex scenarios, guided by traffic priors and optimal control strategies.
The proposed model significantly outperforms traditional platooning methods, such as CACC and CLCC, particularly in safety, efficiency, and robustness. Extensive simulations and HIL experiments validate its effectiveness in handling both static and dynamic disturbances.
Future work will focus on optimizing the model for scalability in larger platoons, refining its performance in real-world settings, and exploring the integration of additional sensory data to enhance situational awareness. These advancements will further advance the reliability and applicability of autonomous platooning systems in commercial deployments.